Mr. Santana is 3 times his daughter Kara's age. Ten years ago, Mr. Santana was 7 times Kara's age.
Which system of equations could you solve to find Mr. Santana's current age? f = father's age and d = daughter's age
A) 3f = d ; f - 10 = 7d - 10
B) f = 3d ; f + 10 = 7(d + 10)
C) f = d/3 ; f + 10 = 7d - 10
D) f = 3d ; f - 10 = 7(d - 10)

Found 2 solutions by fcabanski, addingup:
Answer by fcabanski(1391)

Father (f) is (=) 3 times the daughter's (3d) age. f=3d must be one of the equations. That eliminates A and C.


10 years ago means 10 must be subtracted from things. That eliminates B.


The correct answer is 10. Father's age 10 years ago is f-10 and daughter's age 10 years ago is d-10. He was 7 times her age so f-10=7(d-10).


Answer by addingup(3677)

D is the correct answer.Explain. The problem says:
Father is 3 x daughter, f=3d (D)
------------------------
Next:
10 years ago f= 7d
This was 10 years ago. We deduct 10 from both sides because 10 years ago both f and d were 10 years younger:
f-10= 7(d-10) This is also D)